OpenStudy (anonymous):

Find area of region inside r = 2+2cos(theta) and rsin(theta) = 3/2

OpenStudy (vlery):

if you're looking for theta there's no solution

OpenStudy (catlover5925):

I don't know what type of math u are in but I have never learned this and im a sophmore in an advanced math class...sorry wish I could help

OpenStudy (anonymous):

\[3 /(2\sin \theta) = 2+2\sin \theta\]

OpenStudy (anonymous):

there is one solution 48.6

OpenStudy (anonymous):

The equation \[ r \sin(\theta) =\frac 3 2\\ y=\frac 3 2 \] is a line parallel to the x-axis I do not know what you mean by the are inside a line and a cardioid.

OpenStudy (anonymous):

by the area

OpenStudy (jhannybean):

Area f a region that is bound by polar coordinates: \(A = \dfrac{1}{2}r^2 \theta\) Therefore the area of a olar region is: \(A = \int_a^b \dfrac{1}{2}r^2d\theta\)

OpenStudy (fibonaccichick666):

Since no one has said this, just draw a picture to start. It makes life A LOT easier. I am assuming this is calcI/II and you have just studied integration as pointed out by jhanny. So, start by graphing.